the perimeter of a rectangle is 26 inches. the length of the rectangle is 1.6 times the width of the rectangle. what is the length of the rectangle
Answer by TimothyLamb(4379)

2L + 2w = 26
L = 1.6w
w = (1/1.6)L
w = 0.625L
---
2L + 2w = 26
2L + 2*0.625L = 26
3.25L = 26
L = 26/3.25
L = 8 inches
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---